Net Present Value Method, Present Value Index, and Analysis

Donahue Industries Inc. wishes to evaluate three capital investment projects by using the net present value method. Relevant data related to the projects are summarized as follows:

Product Line
Expansion
Distribution
Facilities
Computer
Network
Amount to be invested $4,000,000 $2,500,000 $500,000
Annual net cash flows:
Year 1 4,200,000 1,000,000 600,000
Year 2 3,600,000 1,200,000 600,000
Year 3 3,000,000 1,300,000 550,000
Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1. Assuming that the desired rate of return is 20%, prepare a net present value analysis for each project. Use the present value of $1 table presented above in your computations. If the net present value is negative, enter a negative amount.

Product Line Expansion Distribution Facilities Computer Network
Present value of net cash flow total $ $ $
Amount to be invested
Net present value $ $ $

2. Determine a present value index for each project. Round your answers to two decimal places.

Present Value Index (Rounded)
Product Line Expansion
Distribution Facilities
Computer Network
Answer #1
Calculation of present value of cash flow
Product line expansion
years Cash flow PVF @20% Present value of cash flow
1 4200000 0.833 3498600
2 3600000 0.694 2498400
3 3000000 0.579 1737000
total present value 7734000
Distribution facilities
years Cash flow PVF @20% Present value of cash flow
1 1000000 0.833 833000
2 1200000 0.694 832800
3 1300000 0.579 752700
total present value 2418500
Computer network
years Cash flow PVF @20% Present value of cash flow
1 600000 0.833 499800
2 600000 0.694 416400
3 550000 0.579 318450
total present value 1234650
Net present value analysis
Product line expansion Distribution facilities Computer network
present value of net cash flow total 7734000 2418500 1234650
amount to be invested -4000000 -2500000 -500000
Net present value 3734000 -81500 734650
2) present value index = present value of cash flow/amount to be invested
present value index (rounded)
Production line expansion 7734000/4000000 = 2.47
Distribution facilities 2418500/2500000 = 0.97
Computer network 1234650/500000   = 2.47

Please give a thumbs up if it is helpful & let me know if any doubt